OpenStudy (elleblythe):

Solve with complete solution: cube root of (4xy^4) sixth root of (12y) + twelfth root of (3x^4y^18) - [3y^2 sixth root of (3x^2)]/sixth root of (y^3) P.S. How do I make the radical sign on here?

OpenStudy (psymon):

You need the equation editor below the chat box to do it. If you want to have the equation stuff in the actual post, you would need to write out the equation in some other question and then copy it in. But Im guessing: \[(\sqrt[3]{4xy ^{4}})(\sqrt[6]{12y})+\sqrt[12]{3x ^{4}y ^{18}}-\frac{( 3y ^{2})\sqrt[6]{3x ^{2}} }{ \sqrt[6]{y ^{3}} }\] Can't tell for sure, though, where the division starts and ends xD

OpenStudy (elleblythe):

@Psymon that's the right equation thanks mate! now would you know how to solve it??